The demand and supply curves for the market are given by: Demand: Qd = 16000 – 24P Supply: Qs = 2000 + 32P A maximum price of $200 is proposed. (e) Does the imposition of the maximum price result in net welfare gains? Explain. (4 marks) (f) Would you advise the implementation of the maximum price? Explain. (3 marks) (g) Suggest an alternate strategy. (2 marks)
Answer #1

(e)

When P = 200,

Qd = 16000 - (24 x 200) = 16000 - 4800 = 11200

Qs = 2000 + (32 x 200) = 2000 + 6400 = 8400

Since consumers can buy only what producers will sell, market quantity traded is 8400. There will be a shortage equal to (Qd - Qs) = (11200 - 8400) = 2800.

The maximum (ceiling price) will generate a net welfare loss because all intended beneficiaries (buyers) cannot purchase the good at the lower ceiling price and a shortage is generated caused by a fall in quantity.

(f)

Since implementation of maximum price will cause a net welfare loss, I will not recommend it.

(g)

Social welfare is maximized when Qd = Qs.

16000 - 24P = 2000 + 32P

56P = 14000

P = 250

Q = 2000 + (32 x 250) = 2000 + 8000 = 10000

This is the welfare-maximizing price-quantity combination, so I will recommend a laissez-faire (free market) strategy.